OpenStudy (anonymous):

medal for the CORRECT answer! Find the value of x in the following rectangle.

OpenStudy (anonymous):

Picture

OpenStudy (anonymous):

that's the picture for this problem

OpenStudy (whpalmer4):

okay, that's a right triangle, right? you know the Pythagorean theorem says that the square of the length of the hypotenuse (the hypotenuse being 5x-4 in this diagram) is the sum of the squares of the other two sides (6, 8). So, that means \[a^2+b^2=c^2\]\[6^2+8^2=(5x-4)^4\] Expand that out and solve for x...

OpenStudy (whpalmer4):

I'll give YOU a medal for a correct answer :-)

OpenStudy (anonymous):

i just need the answer so i can get this overwith.

OpenStudy (anonymous):

it's my last question too.

OpenStudy (jdoe0001):

hehe

OpenStudy (whpalmer4):

No doubt someone else will supply it.

OpenStudy (anonymous):

ugh